# Aging of glass-forming materials following a temperature jump

### v1

@@ Abstract
Aging in a glass forming model in which both trapping effect and delayed response of the free energy landscape (FEL) exist is studied after the temperature is changed. It is confirmed that the trapping effect gives rise to Type-I aging where the relaxation time increases with waiting time regardless of the direction of temperature change, and that the delayed response of the FEL produces Type-II aging where the waiting-time dependence of the relaxation time depends on the direction of temperature change. When both effects exist and the response time of the FEL is appropriate, these effects can be differentiated in the short-time behavior of the temporal relaxation time. It is argued that the material time or the internal clock and the fictive temperature introduced phenomenologically are understood as the concepts describing the delayed response of the FEL to temperature change.
@@ I Introduction
Nonequilibrium systems such as glass-forming materials exhibit properties completely different from those observed in equilibrium systems, where the physical properties observed in experiments after a perturbation is applied depend on the time when the observation is made or started. This phenomenon is called aging [1, 2, 3, 4]. Aging has been observed for many physical properties including enthalpy relaxation [5, 6], dielectric response [7, 8], intermediate scattering function [9] in various kind of systems including spin glasses [10, 11, 12], super-cooled liquids [13], and polymers [14].
Most common experiment is an observation of response to a temperature change. In this experiment, the temperature of heat bath is raised ($T$-up protocol) or lowered ($T$-down protocol) at time $t=0$ and physical properties are measured. The physical properties age and relax to new equilibrium values. Measurement is sometimes made after a waiting time $t_{w}$, and the waiting time dependence of the relaxation is investigated.
@@
Aging has been studied on the basis of a free-energy landscape model by Diezemann [18]. He focused on the difference between translational and rotational relaxation and assumed that the free energy landscape (FEL) responds to the temperature change without delay.
@@
Aging has been phenomenologically understood by considering a material time [19] or an internal clock [8] which explain slow dynamics, where these times pass slower than the actual time. The physical meaning of these concepts are still to be clarified.
In explaining the enthalpy relaxation which shows aging, Tool-Narayanaswamy-Moynihan (TNM) introduced a fictive temperature [21, 22, 23], where the instantaneous structure of a glass in non-equilibrium state below $T_{g}$ is assumed to be the structure of a super-cooled liquid in equilibrium at a fictive temperature. It has been widely used in the analysis of experimental data, allowing quantitative agreement with experimental results on the enthalpy relaxation.
@@ II.1 Model
where $w_{0}$ is an attempt frequency, and $T_{g}$ and $T_{K}$ are the glass transition temperature and the Kauzmann temperature, respectively.
@@ II.2 Calculation of the self-intermediate scattering function
In the FEL approach, the time evolution of a system is represented by a trajectory of the representative point in the FEL, which forms a one dimensional string (reaction coordinate). If we focus on a particular atom and follow its dynamics along one axis in the real space, the dynamics becomes a random walk on the axis. In this paper, we are interested in the relaxation and not in the configurational effects, we consider a random walk on one-dimensional lattice which is described by the master equation [30]
@@
where $P(n,t)$ is the probability of a representative point being in a basin $n$ at a time $t$ and $n$’s constitute a one-dimensional lattice whose lattice constant is $a$.
In this paper, the self-intermediate scattering function (SISF) $F_{s}(k,t)$ is calculated by the following procedure. We first create a random distribution of $\epsilon_{n}$} and assume the temperature of the heat bath is changed from $T_{i}$ to $T_{f}$ at $t=0$. Then, we solve Eq. (9) numerically to obtain $P(n,t)$ under the initial condition that a representative point is on basin $n_{0}$ at $t=0$:
@@
where $t^{\prime}=t-t_{w}$ is the elapsed time from $t=t_{w}$. We discuss aging by $t_{w}$ dependence of the relaxation time of this function. We also define the temporal relaxation time for TTRF, $\tau_{\mathrm{tmp}}(t^{\prime},t_{w})$, by
@@ Appendix B Trapping random walk
$F_{s}(k,t^{\prime},t_{w})=e^{[-2(1-\cos{ka})[\tilde{t}(t^{\prime}+t_{w})-\tilde{t}(t_{w})\}]}.$ (18)

### v2

@@ Abstract
Physical aging is one of the non-equilibrium phenomena where physical properties change over time due to structural relaxation. Aging in spin glass systems has been explained by a trap model on the temperature-independent energy landscape. Meanwhile, in the free energy landscape (FEL) approach to aging phenomena, it is assumed that the FEL responds to temperature changes with a time delay. In this paper, aging in a glass forming model in which both the trapping effect and the delayed response of the FEL exist is studied after the temperature is changed. It is confirmed that the trapping effect gives rise to Type-I aging where the relaxation time increases with waiting time regardless of the direction of temperature change, and that the delayed response of the FEL produces Type-II aging where the waiting-time dependence of the relaxation time depends on the direction of temperature change. When both effects exist and the response time of the FEL is appropriate, these effects can be differentiated in the short-time behavior of the temporal relaxation time. It is argued that the material time or the internal clock and the fictive temperature introduced phenomenologically are understood as the concepts describing the delayed response of the FEL to temperature change.
@@ I Introduction
Nonequilibrium systems such as glass-forming materials exhibit properties completely different from those observed in equilibrium systems, where the physical properties observed in experiments after a perturbation is applied depend on the time when the observation is made or started. This phenomenon is called aging [1, 2, 3, 4]. Aging has been observed for many physical properties including enthalpy relaxation [5, 6, 7, 8], dielectric response [9, 10, 11], intermediate scattering function [12] in various kinds of systems including spin glasses [13, 14, 15], structural glasses [16, 17], and polymers [18, 19].
Most common experiment is an observation of response to a temperature change [20, 21, 22]. In this experiment, the temperature of heat bath is raised ($T$-up protocol) or lowered ($T$-down protocol) at time $t=0$ and physical properties are measured. The physical properties age and relax to new equilibrium values. Measurement is sometimes made after a waiting time $t_{w}$, and the waiting time dependence of the relaxation is investigated.
@@
Aging has been studied on the basis of energy landscape models. The energy landscape approach for glassy systems was first proposed by Stillinger and Weber in the 1980s as the potential energy landscape [26, 27]. Later, energy landscape pictures have extended to an enthalpy landscape [28, 29] and a free energy landscape (FEL) [30]. These approaches succeeded in describing relaxation processes such as a crystal growth rate [29] and a non-linear dielectric response [31]. Diezemann focused on the difference between translational and rotational relaxation and assumed that the FEL responds to the temperature change without delay to explore the aging properties [32].
@@
Aging has been phenomenologically understood by considering a material time [21, 22] or an internal clock [10] which explain slow dynamics, where these times pass slower than the actual time. The physical meaning of these concepts are still to be clarified.
In explaining the enthalpy relaxation which shows aging, Tool-Narayanaswamy-Moynihan (TNM) introduced a fictive temperature [34, 35, 36], where the instantaneous structure of a glass in non-equilibrium state below $T_{g}$ is assumed to be the structure of a super-cooled liquid in equilibrium at a fictive temperature. It has been widely used in the analysis of experimental data, allowing quantitative agreement with experimental results on the enthalpy relaxation [37, 38, 39].
@@ II.1 Model
where $w_{0}$ is an attempt frequency, and $T_{g}$ and $T_{K}$ are the glass transition temperature and the Kauzmann temperature, respectively. A change in $\rho$ derives a change in relaxation dynamics. $\rho=\infty$ corresponds to the simple Debye relaxation [43].
@@
Note that Eqs. (3) and (4) produce the power-law distribution of Eqs. (1) and (2). The jump rate between basins is changed with temperature because the barrier height of FEL changed with temperature, unlike PEL. A model FEL was calculated by using a density functional theory (DFT) [44]. In the framework of DFT, the temperature dependence of FEL is derived from the change in the density distribution.
@@ II.2 Calculation of the self-intermediate scattering function
In the FEL approach, the time evolution of a system is represented by a trajectory of the representative point in the FEL, which forms a one dimensional string (reaction coordinate). If we focus on a particular atom and follow its dynamics along one axis in the real space, the dynamics becomes a random walk on the axis. In this paper, we are interested in the relaxation and not in the configurational effects, we consider a trapping random walk on one-dimensional lattice which is described by the master equation [43]
@@
where $P(n,t)$ is the probability of a representative point being in a basin $n$ at time $t$ and $n$’s constitute a one-dimensional lattice whose lattice constant is $a$.
In this paper, the self-intermediate scattering function (SISF) $F_{s}(k,t)$ is calculated by the following procedure. We first create a random distribution of $\epsilon_{n}$} and assume the temperature of the heat bath is changed from $T_{i}$ to $T_{f}$ at $t=0$. Then, we solve Eq. (9) numerically to obtain $P(n,t)$ under the initial condition that a representative point is in a basin $n_{0}$ at $t=0$:
@@
where $t^{\prime}=t-t_{w}$ is the elapsed time from $t=t_{w}$. We discuss aging by $t_{w}$ dependence of the relaxation time of this function. We also define the temporal relaxation time for TTRF $\tau_{\mathrm{tmp}}(t^{\prime},t_{w})$ by
@@
All variables and parameters are summarized in Table 1. Data and program codes that support the findings of this article are openly available [46].
$W_{n}$ a jump rate of basin $n$ defined by Eq. (7)
$w_{0}$ an attempt frequency
$T_{g}$ the glass transition temperature
$T_{K}$ the Kauzmann temperature
$\epsilon_{n}$ the inherent depth of basin $n$
$T_{\rm{int}}$ the internal temperature defined by Eq. (5)
$T_{i}$ the initial temperature of temperature jump
$T_{f}$ the final temperature of temperature jump
$\tau_{F}$ a relaxation time of the internal temperature
$a$ a lattice constant
$n_{0}$ a basin which a representative point exists at $t=0$
$s$ a sample with a given $\epsilon_{n}$-distribution
$k$ the wave number
$t_{w}$ a waiting time before measurement starts
$t^{\prime}$ an elapsed time after measurement starts
$\tau_{\rm{tmp}}$ the temporal relaxation time for the two-time relaxation function defined by Eq. (15)
Table 1: All variables and parameters
@@ V Discussion
Finally, we comment on future developments of this model. In the present study, we assume a simple exponential function as a delayed response of $T_{\rm{int}}$ (Eq. (8)). It may be modified by calculating heat transfer in amorphous solids. Differences in heat transfer due to differences in local structures may lead to a position-dependent $T_{\rm{}_{int}}$, which may result in some physical behavior such as dynamical heterogeneity [47, 48]. To combine with experimental data, the present model should be extended to describe the aging phenomena including the memory effect [49, 50] and relaxation decoupling [51, 11, 52].
@@ Appendix B Trapping random walk
$F_{s}(k,t^{\prime},t_{w})=e^{-2(1-\cos{ka})[\tilde{t}(t^{\prime}+t_{w})-\tilde{t}(t_{w})]}.$ (18)